A package of 6 fuses are tested where the probability an individual fuse is defective is 0.05 that is 5% of all fuses manufactured are defective . (a) . What is tue probability at least one fuse will be defective ? (b) . What is tue probability that more than one fuse will be defective given that at level one is defective ?
1
Expert's answer
2020-05-19T14:09:03-0400

Binomial distribution problem with "P=0.05" and "n=6"

(a). "P(1\le)"

"P(1\le)=1-P(=0)"

"P(=0) = (1-0.05)^6=0.735092"

"P(1\le)=1-0.735092=0.264908"

(b). "P(1<|1\le)"

"P(>1|\u22651)=P(>1)\/P(\u22651)"

"P(>1)=1-P(=0)-P(=1)=1-0.735-0.232=0.033"

"P(>1|\u22651)=0.033\/0.265=0.124"